NAOJ GW Elog Logbook 3.2

I measured 3 samples of GaAs wafer from LMA.
Reflection measurement is taken at 12°.
Light is not polarized.
Plots show the spectrum of Transmission, Reflection and Absorption ( A = 100% - T - R ).
At 1064nm the absorption is about 4%. After 1300nm absorption becomes negative, down to -2%.
I used the Fresnel laws to determine wether or not the angle 12° would affect the Reflection measurement.
If the light was polarized the difference between 12° and 0° would be 2.5%
But the spectrophotometer SolidSpec-3700 have unpolarized light, and the difference between 12° and 0° is a factor of 1.5e-5 (0.0015%).
Not enough to affect the measurement, unless the light was polarized for some reason.

I made cross-checkings for Shoda-san's measuremets of "SPARE MIRROR" at Kagra site.
1) OSEM flag position on mirror
Within the accuracy of 1mm, these measurements are almost consistent with the design.
Only a measured value of the upper edge of the mirror was wrong.
2) Relative position of OSEM flag respect to the through-hole
After the hanging, it seems that the mirror is placed on 6mm upside from the mirror box.
I confirm it by making a drawing as attached.

The Remaining 37 OSEMs are now being cured in the ISO-1 clean booth in the ATC multipurpose experimental room!
Don't touch them for a couple of days.
Previously I did this 30 OSEM coils, and today 37coils, that means there will be 67 OSEMs in total. The number is sufficient for iKAGRA, but not for bKAGRA, as at least 70 OSEMs will be required for the bKAGRA (PRM, PR2, PR3, BS, SRM, SR2, SR3).
There are some OSEM coils washed or unwashed and used for trials of several process for the assembly, and they should be added to the number to fulfill the required numbers.

I draw on Inventor the mirror mount for the tama-sized mirrors (d100 x 60).
The picture shows the transation stage part where the mount have to be fixed.
I attach a pdf of the drawing (actually, the top view shouldn't be red, it's a problem of the conversion to pdf, so I also attach the bitmap of the topview).
The yellow part is peek plastic, which is in contact with the mirror.
The red part is the part that have to be fixed permanently to the translation stage.
I put a precision pin, to attach the mirror mount always in the same position.

I finished the (final?) simulations regarding the scattering of the BS recoil mass with the revised model of the BS, which I got from Fabian.
Basically, as expected, no significant effect could be found on KAGRAs sensitivity.
However, I have widen the previous model with the view on the clear-aperture of 195mm. Additionally, I found a way how to add a quasi-elliptical light-source for the mirrors surface, which is of great advantage for increasing the resolution of the results and to minimize the statistical error.
Right now, the error for the maximum values is ca. 10% but further decreasing it is no problem because of the mentioned reasons. Maybe, I will make another simulation run on Monday with increased resolution, also because of having better results for my paper.

2nd IM was almost assembled and shipped to Kamioka together with the 1st IM.
In the 2nd IM, one of the screw was stucked when assembling the side wall (spring side).
I think shorter screw can be inserted, but the long screw cannot be screwed fully now.
Anyway, it works without that screw.
And the heat shrink tube is running out.
So I did not connected the top picomotor cable.

Plus 5 OSEM coil bodies are washed (by ethanol), and then 67+5 = 72 OSEM coil bodies will be available in the end.

A short notice: I confirmed the LOCTITE vac seal can survive with ethanol in the ultrasonic bath.

I took some picture while the workers were repairing and repainting the floor during these three days.
Their work was basically
- Scratch the old floor and clean to the concrete.
- Paint the concrete with a black paint.
- Put a layer of green paint
- Put a second layer of green paint.

Ikenoue, Fabian, Fujii, Shoda,
Ikenoue-san told us how to use the winch system.
We tested the winching with 0.2mm piano wire and 0.65mm Tungsten wire.
We successfully lifted 4 kg and 7.2 kg weights by both wires.
The new winch system is fine for thick wires such as 0.65mm Tungsten wires.
We should add is some marks to indicate the way to turn the nobs.
And when you put the winch system on the IM, please remove the cable clamp and put the cable inside the IM, otherwise the cable will be sandwitched.

37 OSEM coil bodies are now packed into clean bags. (Akutsu, Fujii)

I tested how He would influence the time of the cooling-down and the warming-up of the cryostat.
For testing the cooling-down, I injected 20Pa of He gas in the tank and started the cryo-pumps. Interestingly, by reaching 250K, the temperature curve started to decrease less than the curve for pure vacuum. It then approaches to a temperature of ca. 180K. By pumping off the He inside the tank, the curve decreases as usual again (first picture, red curve). It seems that the isolation around the cold finger is not very good and He started to become a heat-bridge...
Warming-up the cryostat, on the other side, becomes now a very fast thing by injecting He gas: after 50h, I reached room temperature again. This is very convenient compared to the one week (!) we needed with pure vacuum.
The temperature curve for warming-up the cryostat can be seen in the second attached picture.
(Because of the low-temperatures, when I injected some He into the tank, I could not control the final pressure properly. That is why I got 62Pa in the end and not 20Pa as I was aimed for...)

Mitsui-san teached me how to use the spectrophotometer SolidSpec-3700 in the optic shop at ATC.
I measured reflection and transmission at 1064nm of the GaAs sample: S-I axt 8000010699 71
Sample thickness 0.4mm
Reflection = 46.77% (at 12°)
Transmission = 51.16%
Absorption = 100-R-T = 2.07%

Fujii, Hirata, Shoda
We assembled the IM for PR3 except for the wire clamps and the OSEM flags.
Please see the pictures.
The reason why we did not attach the wire clamps is the pins do not fit well. We can fix it by pushing, but if we do that, it seems difficult to remove the clamp base.
The clamp bases and clamps are paired so that the pins for alignment fit well.
But the side board, where the clamp base is attached, is not paired while the clamp base also has alignment pins.
Also, the picomotor length did not fit to the cad drawing.
I could not find why since we do not have cad file it self at that time.
We just adjust the length of the picomotor so that the the position of the moving mass is almost at the center (by eyes).
Note: one of the picomotor driver did not work well. It goes on forward, but not on backward.

I have measured the surface-map of two SiC samples (both polished) with Zygos NewView instrument.
One sample is from Kyocera and the other one from a cheaper company. For both samples it is actually quite difficult to measure the rms roughness as in both samples, the surface is featured by a lot of holes, partly having depths of 2-6 mum. Therefore, I got rms values ranging from 1 to 10 nm for both samples. Nevertheless, from the maps itself one can get the conclusion that kyoceras polishing (the first picture) is better as, apart from the holes, it looks smoother.

In order to obtain an estimation of the correction factor between bulk absorption and surface absorption on the same material (GaAs), I made some calculations.
The assumptions are a bit strong:
- the signal from the photodiode detector is proportional to the phase change of the probe beam.
- the phase change of the probe beam is proportional to the optical path change.
- the optical path change is proportional to the integral of temperature along the beam path inside the material.
I calculate the temperature distribution inside a GaAs of thickness 400micron, absorption of 12ppm on a surface of 10micron (Figure1) and absorption of 12ppm inside all the bulk 400micron (Figure2).
Calculate the integral of the temperature in dz (depth), as a function of r (radius) for both bulk and surface absorption (Figure3). Optical path change.
Divide the optical path change of bulk absorption by the optical path change of surface absorption. Plot the ratio as a function of r (Figure4).
I plot the same thing in a wider range, up to 5cm or radius to see how the graph goes to zero.
In the second figure I take the second derivative of the optical path with respect to the radius. The second derivative should de proportional to the focusing power, hence, to the signal of the absorption system.
It is the same for the bulk and for the surface absorption

I continued the IM assembly.
1. Attach the pin connector to the picomotors.
2. Move the picomotors and adjust its position.
3. Attach the picomotors to the base plate.
4. attach the spring and the side wall of IM box.
5. same for the upper picomotor.
*What I noticed:
- The spring is compressed a lot. So the spring winds even we have the pole inside. (the gap between the spring and the pole is not small.)
- the tube for the connection pins are too loose to be fixed at their position. At this time, I fixed the tube and the cable using the cable clamps so that the tubes do not move.
But if we use the same tube for the other parts, we need to shrink these using a heat gun or something.
The procedure remained is connecting upper board to the base(, which needs more than two people).

The winch systems for IMs are ready in the ATC clean booth.

Members: Tatsumi, Manuel.
Big translation stage have been delivered. The company man came last thursday and today. We temporary mounted the parts on a small optical table, we configured the controllers using its software. It works properly.
Next steps before setting the big translation stage are:
- To measure the absorption of the two Shinkousha sapphire samples that Hirose-san kindly sent to us.
- To make some power measurements of the pump (after and before the chopper), in order to see how the power measurement depends on the modulation frequency.
- To understand better the beam profile of the probe beam in order to be able to make a detail design of how to move the parts of the absorption system and make enough space to place the big mirrors.

Members: Tatsumi, Manuel
Yesterday we moved many things in Tama central room from downstairs to upstairs. Picture1.jpg
Today a company came to move the optical tables. They moved the absorption measurement system optical table is inside the booth beside the stairs. Picture2.jpg
Now the floor is free (exept of the green shelves we will move soon) and ready to be repainted. Picture3.jpg